Ncert Solutions Maths class 12th

Get insights from 2.5k questions on Ncert Solutions Maths class 12th, answered by students, alumni, and experts. You may also ask and answer any question you like about Ncert Solutions Maths class 12th

  • 2.5k

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 65

    Followers

Have a question about Ncert Solutions Maths class 12th? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Probability  of  drawing  2  green  and  1  blue  balls                 =P(G).P(G).P(B)+P(G).P(B).P(G)+P(B).P(G).P(G)                 =38.27.26+38.27.26+28.37.26=12336+12336+12336=36336=328Hence,  the  correct  option  is  (a).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Let  E1  be  the  event  of  getting  even  number  on  the  die.  E2  be  the  event  of  selecting  a  spade  card.∴    P(E1)=36=12  and  P(E2)=1352=14So,           P(E1∩E2)=P(E1).P(E2)=12.14=18Hence,  the  correct  option  is  (c).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Let  G  denotes  the  girl  and  B  denotes  the  boy  of  the  given  family.So,  n(S)={(BGG),(GBG),(GGB),(GBB),(BGB),(BBG),(BBB),(GGG)}Let  E1  be  the  event  that  the  family  has  atleast  one  girl∴            E1={(BGG),(GBG),(GGB),(GBB),(BGB),(BBG),(GGG)}.  E2  be  the  event  that  the  eldest  child  is  a  girl.∴            E2={(GBG),(GGB),(GBB),(GGG)}(E1∩E2)={(GBB),(GGB),(GBG),(GGG)}∴P(E2/E1)=P(E1∩E2)P(E1)=4/87/8=47Hence,  the  correct  option  is  (d).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Given  that:  P(A)=0.4,  P(B)=0.3,  P(C)=0.2Also,  P(A¯)=1−0.4=0.6,  P(B¯)=1−0.3=0.7,  P(C¯)=1−0.2=0.8∴  Probabilities of  two  hits                 =P(A).P(B).P(C¯)+P(A).P(B¯).P(C)+P(A¯).P(B).P(C)                 =0.4×0.3×0.8+0.4×0.7×0.2+0.6×0.3×0.2                 =0.096+0.056+0.036=0.188Hence,  the  correct  option  is  (b).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

According  to  Question  74,Let  E1  be  event  that  first  ball  is  red.          E2 be  event  that  exactly  two  of  the  three  balls  are  red.∴  P(E1)=P(R).P(R).P(B)+P(R).P(R).P(R)+P(R).P(B).P(R)+P(R).P(B).P(B)                   =58.47.36+58.47.36+58.37.46+58.37.26                   =60336+60336+60336+30336=210336P(E1∩E2)=P(R).P(B).P(R)+P(R).P(R).P(B)                          =58.37.46+58.47.36=60336+60336=120336∴P(E2/E1)=P(E1∩E2)P(E1)=120/336210/336=47Hence,  the  correct  option  is  (b).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Given  that:Bag  contains  5  red  and  3  blue  balls.  of  getting  exactly  one  red  ball  if  3  balls  are  randomly  drawn  without  replacement.P(R).P(B).P(B)+P(B).P(R).P(B)+P(B).P(B).P(R)  =58.37.26+38.57.26+38.27.56=30336+30336+30336=90336=1556Hence,  the  correct  option  is  (c).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

Given  that:E  and  F  are  two  independent  events  such  thatP(E)=0.3,  and  P(E∪F)=0.5∴        P(E∪F)=P(E)+P(F)−P(E∩F)                             0.5=0.3+P(F)−P(E).P(F)⇒             0.5−0.3=P(F)[1−P(E)]      ⇒0.2=P(F)(1−0.3)⇒                        0.2=P(F).(0.7)∴                        P(F)=0.20.7=27Now  P(E/F)−P(F/E)=P(E∩F)P(F).P(E∩F)P(E)        =P(E).P(F)P(F).P(E).P(F)P(E)=P(E)−P(F)=310−27=170Hence,  the  correct  option  is  (c).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

  A  and  B  are  two  independent  events∴          P (A∩B)=P (A).P (B)Hence,   the  correct  option  is   (c).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Objective Type Questions as classified in NCERT Exemplar

Sol:

G i v e n     t h a t : P ( A ) = 3 8 ,     P ( B ) = 5 8     a n d     P ( A ∪ B ) = 3 4 ∴                 P ( A ∪ B ) = P ( A ) + P ( B ) − P ( A ∩ B )                                                           3 4 = 3 8 + 5 8 − P ( A ∩ B ) ⇒               P ( A ∩ B ) = 3 8 + 5 8 − 3 4 = 1 4 N o w     P ( A / B ) . P ( A ' / B ) = P ( A ∩ B ) P ( B ) . P ( A ' ∩ B ) P ( B )                                                                                                         = P ( A ∩ B ) P ( B ) . P ( B ) − P ( A ∩ B ) P ( B )                                                                                                         = 1 4 5 . ( 5 8 − 1 4 ) 5 = 2 5 . 3 5 = 6 2 5 H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( d ) .

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago